The lowest pressure attainable using the best available vacuum techniques is about 10−12N/m2. At such a pressure, how many molecules are there per cm3 at 3 ∘C?
Answer #1

Use ideal gas equation to find the required number of molecules after finding the number of moles as shown belowfrom Ideal gas eqy - PV = nRT {v=1cm3 = 100m =) (10-12 N/m2) (10-6m3) = nR (273+3)K => m= 4.36X 10-22 mole I mole = 6:022x102***********************************************************************************************
This concludes the answers. If there is any mistake or omission, let me know immediately and I will fix it....